### v1

@@ 1 Introduction
The risk model to be discussed in this paper is given below. Let $X=\{X_{t},t\geq 0\}$ be a SNLP with characteristic triplet $(\gamma,\sigma,\nu)$ defined on the filtered probability space $(\Omega,\mathcal{F},\mathbb{F}=\{\mathcal{F}_{t},t\geq 0\},P)$, where $\sigma\geq 0$ is the continuous diffusion coefficient, $\gamma\in\mathbb{R}$ is the linear drift coefficient and $\nu$ is the Lévy measure that satisfies $\nu(-\infty,0)=0$ and $\int_{0}^{\infty}(1\wedge z^{2})\nu(\mathrm{d}z)<\infty$. The law of $X$ is denote by $P_{x}$ when it starts at $x\in\mathbb{R}$, and for convenience, $P$ is used instead of $P_{0}$. The associated expectation operators are denoted as $E_{x}$ and $E$, respectively. The refracted Lévy process $R=\{R_{t}:t\geq 0\}$ is described by
$R_{t}=X_{t}-\delta\int_{0}^{t}\mathbf{1}\{R_{s}\geq b\}\mathrm{d}s,\ \ \ \ t\geq 0,\ \ \delta\geq 0,$
where $b\geq 0$ is a threshold level representing the insurer’s solvency capital requirement, $\delta\geq 0$ is a refraction parameter and $\mathbf{1}\{A\}$ is the indicator function of event $A$. Adapting the premium in this way can be seen as a modern investment strategy. Specifically, when the insurance company’s financial situation is regular or good, meaning its surplus is at or above level $b$, it invests at rate $\delta$ and the surplus process follows the risk process $Y=\{Y_{t}=X_{t}-\delta t,t\geq 0\}$. When the company is in financial distress, the risk process is $X=\{X_{t},t\geq 0\}$ without any investment. Equivalently, $R_{t}=Y_{t}+\delta\int_{0}^{t}\mathbf{1}\{R_{s}<b\}\mathrm{d}s$, where $Y$ is regarded as the underlying risk process during normal business periods, and a restructuring is undertaken once $R$ drops below $b$, in the sense that the premium applied to large claims is temporarily increased.
@@ 4 Optimal impulse control strategy
We introduce a candidate for the optimal strategy in the dividend problem (1.3), known as the impulse control strategy $\pi_{(c_{1},c_{2})}$, which represents an important type of strategy for impulse control problem. Specifically, we set two levels $c_{1}$ and $c_{2}$ such that $c_{1}\geq 0$ and $c_{2}>c_{1}+\beta$, and fix a set of the stopping times $\{\theta_{k}^{(c_{1},c_{2})},k=1,2,\cdots\}$, where $\theta_{k}^{(c_{1},c_{2})}$ is given by, for $k=1,2,\cdots$,
$\theta_{k}^{(c_{1},c_{2})}:=\inf\{t\geq 0:R_{t}>(R_{0}\vee c_{2})+(c_{2}-c_{1})(k-1)\}.$
The impulse control strategy $\pi_{(c_{1},c_{2})}=\{D_{t}^{(c_{1},c_{2})},t\geq 0\}$ is defined as
$D_{t}^{(c_{1},c_{2})}:=\mathbf{1}\{\theta_{1}^{(c_{1},c_{2})}<t\}\big{(}(R_{0}\vee c_{2})-c_{1}\big{)}+\sum_{k=2}^{\infty}\mathbf{1}\{\theta_{k}^{(c_{1},c_{2})}<t\}(c_{2}-c_{1}).$
The controlled process $U^{\pi}$ becomes $U_{t}^{(c_{1},c_{2})}=R_{t}-D_{t}^{(c_{1},c_{2})}$. Note that in terms of $U_{t}^{(c_{1},c_{2})}$, $\theta_{k}^{(c_{1},c_{2})}$ can be regarded as the time when the process $U^{(c_{1},c_{2})}$ exceeds $c_{2}$ for the $k$-th time, i.e. $\theta_{1}^{(c_{1},c_{2})}:=\inf\{t\geq 0:U_{t}^{(c_{1},c_{2})}>c_{2}\}$ and $\theta_{k}^{(c_{1},c_{2})}:=\inf\{t>\theta_{k-1}^{(c_{1},c_{2})}:U_{t}^{(c_{1},c_{2})}>c_{2}\}$ for $k\geq 2$. The expected discounted dividend function and ruin time for the strategy $\pi_{(c_{1},c_{2})}$ are denoted as $V_{(c_{1},c_{2})}$ and $T^{(c_{1},c_{2})}$, respectively. In summary, the impulse strategy $\pi_{(c_{1},c_{2})}$ is to reduce the risk process $U^{(c_{1},c_{2})}$ to $c_{1}$ whenever it exceeds level $c_{2}$. Here, $c_{2}-c_{1}>\beta$ ensures that shareholders receive dividends after paying the transaction costs, while $c_{1}\geq 0$ is based on admissible constraint (1.2).
@@ 4.1 Expected discounted dividend function for impulse strategy $\pi_{(c_{1},c_{2})}$
Note that $V_{(c_{1}^{*},c_{2}^{*})}$ is an increasing function. In fact, $V_{c_{2}^{*}}:=V_{(c_{1}^{*},c_{2}^{*})}$ is the expected discounted dividend function associated with the barrier strategy at level $c_{2}^{*}$ for the de Finetti’s dividend problem with exponential Parisian ruin and an ultimate bankrupt barrier.
@@ 4.2 Optimality of impulse strategy $\pi_{(c_{1}^{*},c_{2}^{*})}$
We first prove that $V_{(c_{1}^{*},c_{2}^{*})}$ satisfies the HJB inequality (3.4). For $l\leq x<c_{2}^{*}$, to verify that $(\mathcal{A}-q-m\mathbf{1}\{x<0\})V_{(c_{1}^{*},c_{2}^{*})}(x)=0$, it is sufficient, by Proposition 4.1, to show that the process $\{e^{-q(t\wedge\tau)}\vartheta^{(q+m,q)}(U^{(c_{1}^{*},c_{2}^{*})}_{t\wedge\tau},-l)\}_{t\geq 0}$ is a $P_{x}$-martingale, where $\tau:=T\wedge k_{c_{2}^{*}}^{+}$. Note that no dividends are paid before time $t\wedge\tau$, i.e. $U^{(c_{1}^{*},c_{2}^{*})}_{t\wedge\tau}=R_{t\wedge\tau}$ for all $t\geq 0$. Let $\tilde{\tau}:=(\tau-t)^{+}=\max\{0,\tau-t\}$ and $\tilde{R}_{\tilde{\tau}}:=R_{t+\tilde{\tau}}$. Using the strong Markov property and (3.1), together with the fact that $\vartheta^{(q+m,q)}(R_{\tau},-l)=\vartheta^{(q+m,q)}(c_{2}^{*},-l)\mathbf{1}\{k_{c_{2}^{*}}^{+}<T\}$, we have,
$E_{x}[e^{-q\tau}\vartheta^{(q+m,q)}(R_{\tau},-l)\mid\mathcal{F}_{t}]$
$=\mathbf{1}\{t\leq\tau\}e^{-qt}E_{x}[e^{-q(\tau-t)}\vartheta^{(q+m,q)}(R_{\tau},-l)\mid\mathcal{F}_{t}]+\mathbf{1}\{\tau<t\}e^{-q\tau}\vartheta^{(q+m,q)}(R_{\tau},-l)$
$=\mathbf{1}\{t\leq\tau\}e^{-qt}E_{R_{t}}[e^{-q\tilde{\tau}}\vartheta^{(q+m,q)}(\tilde{R}_{\tilde{\tau}},-l)]+\mathbf{1}\{\tau<t\}e^{-q\tau}\vartheta^{(q+m,q)}(R_{\tau},-l)$
$=\mathbf{1}\{t\leq\tau\}e^{-qt}\vartheta^{(q+m,q)}(R_{t},-l)+\mathbf{1}\{\tau<t\}e^{-q\tau}\vartheta^{(q+m,q)}(R_{\tau},-l)$
$=e^{-q(t\wedge\tau)}\vartheta^{(q+m,q)}(R_{t\wedge\tau},-l),$
which implies that the process $\{e^{-q(t\wedge\tau)}\vartheta^{(q+m,q)}(U^{(c_{1}^{*},c_{2}^{*})}_{t\wedge\tau},-l)\}_{t\geq 0}$ is indeed a $P_{x}$-martingale.
Alternatively, a more general verification of (3.4) can be established as follows. As derived from Proposition 2.2, the derivative of $\vartheta^{(q+m,q)}(\cdot,-l)$ fails to exist at 0 or $b$ when $X$ is of BV, and the second left derivative of $\vartheta^{(q+m,q)}(\cdot,-l)$ at $c_{2}^{*}$ does not necessarily equal zero when $X$ is of UBV. Accordingly, $(\mathcal{A}-q-m\mathbf{1}\{x<0\})V_{(c_{1}^{*},c_{2}^{*})}(x)$ is not well-defined. Therefore, we moderate our claim to demonstrate that the following result holds for any $t\in[0,T]$,
$\int_{0}^{t}e^{-qs-L^{\pi}(s)}(\mathcal{A}-q-m\mathbf{1}\{\tilde{U}_{s}^{(c_{1}^{*},c_{2}^{*})}<0\})V_{(c_{1}^{*},c_{2}^{*})}(\tilde{U}_{s}^{(c_{1}^{*},c_{2}^{*})})\mathrm{d}s\leq 0,\ \ \ a.s.,$ (4.13)
@@ 5 Examples
The existence and satisfiability of the optimal impulse control $\pi_{(c_{1}^{*},c_{2}^{*})}$ are showed in Proposition 4.2. Combined with Theorem (4.5), this result enables the development of a method for numerically computing the optimal thresholds. Compared to the de Finetti’s problem discussed in Loeffen[26], the additional complexity involves characterizing the optimal values of parameters $c_{1}$ and $c_{2}$, which forms a two-dimensional optimization problem, in contrast to the one-dimensional problem of finding the optimal barrier strategy. In this section, we analyze two examples for which it is able to present the numerical results of the optimal points $(c_{1}^{*},c_{2}^{*})$, including refracted Brownian risk model and refracted Cramér-Lundberg model with exponential claims.

### v2

@@ 1 Introduction
The risk model to be discussed in this paper is given below. Let $X=\{X_{t},t\geq 0\}$ be a SNLP with characteristic triplet $(\gamma,\sigma,\nu)$ defined on the filtered probability space $(\Omega,\mathcal{F},\mathbb{F}=\{\mathcal{F}_{t},t\geq 0\},P)$, where $\sigma\geq 0$ is the diffusion coefficient, $\gamma\in\mathbb{R}$ is the drift coefficient and $\nu$ is the Lévy measure that satisfies $\nu(-\infty,0)=0$ and $\int_{0}^{\infty}(1\wedge z^{2})\nu(\mathrm{d}z)<\infty$. The law of $X$ is denote by $P_{x}$ when it starts at $x\in\mathbb{R}$, and for convenience, $P$ is used instead of $P_{0}$. The associated expectation operators are denoted as $E_{x}$ and $E$, respectively. The refracted Lévy process $R=\{R_{t}:t\geq 0\}$ is described by
$R_{t}=X_{t}-\delta\int_{0}^{t}\mathbf{1}\{R_{s}\geq b\}\mathrm{d}s,\ \ \ \ t\geq 0,\ \ \delta\geq 0,$ (1.1)
where $b\geq 0$ is a threshold level, $\delta\geq 0$ is a refraction parameter and $\mathbf{1}\{A\}$ is the indicator function of event $A$. Economically, the threshold $b\geq 0$ represents a solvency or capital management level (e.g., a regulatory intervention threshold) at which the insurer switches between operating regimes, while the refraction parameter $\delta\geq 0$ quantifies the strength of the adjustment made in the high-surplus regime. When $R_{t}\geq b$, the company is viewed as well capitalised and the surplus is modified at rate $\delta$ (e.g. through active investment, intensified risk mitigation, or an additional dividend or reinsurance outflow), whereas for $R_{t}<b$ the surplus follows the baseline risk process $Y=\{Y_{t}=X_{t}-\delta t,t\geq 0\}$ without such adjustments. Thus $b$ encodes the trigger at which capital management actions are activated, and $\delta$ measures the intensity of these actions, allowing the model to capture a range of investment, dividend, and restructuring policies within a unified framework.
@@ 3.1 Exit problem
On the basis of (3.1), we now derive several limiting cases. A detailed proof of Corollary 3.2 is given in Appendix A.3, the other corollaries follow by similar arguments.
Corollary 3.2.
Letting $l\rightarrow\infty$, ruin is governed solely by the Parisian exponential delay, i.e. $T=\tau_{p}$. For $x\in(-\infty,c)$, we have
$E_{x}[e^{-qk_{c}^{+}}I\{k_{c}^{+}<\tau_{p}\}]=\frac{F^{(q+m,q)}(x)}{F^{(q+m,q)}(c)},$ (3.2)
where $H^{(q+m,q)}(x):=e^{\phi(q+m)x}\big(1-m\int_{0}^{x}e^{-\phi(q+m)(x-y)}W^{(q)}(x-y)\mathrm{d}y\big)$ and
$F^{(q+m,q)}(x):=H^{(q+m,q)}(x)+\delta\mathbf{1}{\{x\geq b\}}$
$\cdot\int_{b}^{x}\mathbb{W}^{(q)}(x-y)\Big(\big(\phi(q+m)-mW^{(q)}(0)\big)e^{\phi(q+m)y}-m\int_{0}^{y}e^{\phi(q+m)z}W^{(q)}{{}^{\prime}}(y-z)\mathrm{d}z\Big)\mathrm{d}y.$
Corollary 3.3.
For $m=0$, the model (1.1) is subject to no restriction on the time it may spend below zero, and ruin is declared at the first down-crossing of the lower bankruptcy barrier l-l, i.e. $T=k_{-l}^{-}$. For $x\in[-l,c)$, we have
$E_{x}\bigl[e^{-qk_{c}^{+}}\mathbf{1}{\{k_{c}^{+}<k_{-l}^{-}\}}\bigr]=\frac{w^{(q)}(x;-l)}{w^{(q)}(c;-l)},$
which is consistent with (2.5). In particular, as $l\to 0$ (equivalently, $m\to\infty$), ruin is defined in the classical sense as the first passage time below zero, i.e. $T=k_{0}^{-}$.
Corollary 3.4.
For $\delta=0$, the model (1.1) reduces to SNLP $X$. Then, for $x\in[-l,c)$,
$E_{x}[e^{-q\tau_{c}^{+}}\mathbf{1}\{\tau_{c}^{+}<T\}]=\frac{g^{(q+m,q)}(x,l)}{g^{(q+m,q)}(c,l)},$ (3.3)
which is consistent with Proposition 3.1 of Frostig and Keren-Pinhasik[14]. Further, letting $l\rightarrow\infty$, we have $T=\tau_{p}$, and for $x\in(-\infty,c)$,
$E_{x}[e^{-q\tau_{c}^{+}}\mathbf{1}\{\tau_{c}^{+}<{\tau}_{p}\}]=\frac{H^{(q+m,q)}(x)}{H^{(q+m,q)}(c)},$
which recovers the result in Baurdoux, Pardo, Pérez and Renaud[4] and Lemma 2.1 of Landriault, Renaud, and Zhou[21]. Moreover, for $m=0$, we have $T=\tau_{-l}^{-}$, and (3.3) coincides with (2.3).
@@ 4 Optimal impulse control strategy
We introduce a candidate for the optimal strategy in the dividend problem (1.4), known as the impulse control strategy $\pi_{(c_{1},c_{2})}$, which represents an important type of strategy for impulse control problem. Specifically, we set two levels $c_{1}$ and $c_{2}$ such that $c_{1}\geq 0$ and $c_{2}>c_{1}+\beta$, and fix a set of the stopping times $\{\theta_{k}^{(c_{1},c_{2})},k=1,2,\cdots\}$, where $\theta_{k}^{(c_{1},c_{2})}$ denotes the time at which the process $U^{(c_{1},c_{2})}$ exceeds $c_{2}$ for the $k$-th time, i.e. $\theta_{1}^{(c_{1},c_{2})}:=\inf\{t\geq 0:U_{t}^{(c_{1},c_{2})}>c_{2}\}$ and $\theta_{k}^{(c_{1},c_{2})}:=\inf\{t>\theta_{k-1}^{(c_{1},c_{2})}:U_{t}^{(c_{1},c_{2})}>c_{2}\}$ for $k\geq 2$. The expected discounted dividend function and ruin time for the strategy $\pi_{(c_{1},c_{2})}$ are denoted as $V_{(c_{1},c_{2})}$ and $T^{(c_{1},c_{2})}$, respectively. We also denote $\tau_{p}^{\pi}$ and $k_{-l}^{\pi}$ by $\tau_{p}^{(c_{1},c_{2})}$ and $k_{-l}^{(c_{1},c_{2})}$, respectively. In summary, the impulse strategy $\pi_{(c_{1},c_{2})}$ is to reduce the risk process $U^{(c_{1},c_{2})}$ to $c_{1}$ whenever it exceeds level $c_{2}$. Here, $c_{2}-c_{1}>\beta$ ensures that shareholders receive dividends after paying the transaction costs, while $c_{1}\geq 0$ is based on admissible constraint (1.3).
@@ 4.1 Expected discounted dividend function for impulse strategy $\pi_{(c_{1},c_{2})}$
Note that $V_{(c_{1}^{*},c_{2}^{*})}$ is an increasing function. In fact, $V_{c_{2}^{*}}:=V_{(c_{1}^{*},c_{2}^{*})}$ is the expected discounted dividend function associated with the barrier strategy at level $c_{2}^{*}$ for the de Finetti’s dividend problem with exponential Parisian ruin and ultimate bankrupt barrier.
Remark 4.4.
In certain limiting parameter regimes, the optimal dividend strategy $\pi_{(c_{1}^{*},c_{2}^{*})}$ and its value function $V_{(c_{1}^{*},c_{2}^{*})}(x)$ reduce to those of several well-studied dividend problems. (i)As $l\to\infty$, we have $T^{(c_{1}^{*},c_{2}^{*})}=\tau_{p}^{(c_{1}^{*},c_{2}^{*})}$, and $V_{(c_{1}^{*},c_{2}^{*})}$ can be expressed in terms of $F^{(q+m,q)}$, since $\lim_{l\to\infty}\vartheta^{(q+m,q)}(x,-l)=F^{(q+m,q)}(x)$ by Corollary 3.2. If in addition $\delta\to 0$, then problem (1.4) reduces to SNLP dividend problem with Parisian exponential delay and transaction costs, and $V_{(c_{1}^{*},c_{2}^{*})}$ formally coincides with that in Theorem 4.1 of Renaud[30], since $\lim_{l\to\infty,\delta\to 0}\vartheta^{(q+m,q)}(x,-l)=H^{(q+m,q)}(x)$ by Corollary 3.4.
(ii) As $l\to 0$ (or $m\to\infty$), we have $T^{(c_{1}^{*},c_{2}^{*})}=k_{0}^{(c_{1}^{*},c_{2}^{*})}$ (classic ruin), and $V_{(c_{1}^{*},c_{2}^{*})}$ can be expressed in terms of $w^{(q)}(x;0)$, since $\lim_{l\to 0}\vartheta^{(q+m,q)}(x,-l)=w^{(q)}(x;0)$ by Corollary 3.3. If in addition $\delta\to 0$, then problem (1.4) reduces to SNLP dividend problem with classic ruin and transaction costs, and $V_{(c_{1}^{*},c_{2}^{*})}$ coincides with that in Corollary 4 of Loeffen[27], since $\lim_{l\to 0,\delta\to 0}\vartheta^{(q+m,q)}(x,-l)=W^{(q)}(x)$ by Corollary 3.4. Moreover, when $\delta,\beta\to 0$, problem (1.4) further reduces to SNLP dividend problem with classic ruin and without transaction costs, and $V_{(c_{1}^{*},c_{2}^{*})}$ matches the one in Proposition 1 of Loeffen[26], in particular, $(c_{1}^{*},c_{2}^{*})$ collapses to a single optimal barrier $b^{*}$.
@@ 4.2 Optimality of impulse strategy $\pi_{(c_{1}^{*},c_{2}^{*})}$
We first prove that $V_{(c_{1}^{*},c_{2}^{*})}$ satisfies the HJB inequality (3.6). For $l\leq x<c_{2}^{*}$, to verify that $(\mathcal{A}-q-m\mathbf{1}\{x<0\})V_{(c_{1}^{*},c_{2}^{*})}(x)=0$, it is sufficient, by Proposition 4.1, to show that the process $\{e^{-q(t\wedge\tau)-L^{(c_{1}^{*},c_{2}^{*})}(t\wedge\tau)}\vartheta^{(q+m,q)}(U^{(c_{1}^{*},c_{2}^{*})}_{t\wedge\tau},-l)\}_{t\geq 0}$ is a $P_{x}$-martingale, where $\tau:=T^{(c_{1}^{*},c_{2}^{*})}\wedge\hat{k}_{c_{2}^{*}}^{+}$ with $\hat{k}_{c_{2}^{*}}^{+}$ denoting the first time that $U^{(c_{1}^{*},c_{2}^{*})}$ exceeds the level $c_{2}^{*}$, and $L^{(c_{1}^{*},c_{2}^{*})}(t):=\int_{0}^{t}m\mathbf{1}\{U_{s}^{(c_{1}^{*},c_{2}^{*})}<0\}\mathrm{d}s$. Note that no dividends are paid before time $t\wedge\tau$, i.e. $\tau=T\wedge k_{c_{2}^{*}}^{+}$, $U^{(c_{1}^{*},c_{2}^{*})}_{t\wedge\tau}=R_{t\wedge\tau}$ and $L^{(c_{1}^{*},c_{2}^{*})}(t)=L(t):=\int_{0}^{t}m\mathbf{1}\{R_{s}<0\}\mathrm{d}s$ for all $t\geq 0$. Let $\tilde{\tau}:=(\tau-t)^{+}=\max\{0,\tau-t\}$ and $\tilde{R}_{\tilde{\tau}}:=R_{t+\tilde{\tau}}$. By strong Markov property and (3.1), together with the fact $\vartheta^{(q+m,q)}(R_{\tau},-l)=\vartheta^{(q+m,q)}(c_{2}^{*},-l)\mathbf{1}\{k_{c_{2}^{*}}^{+}<T\}$, we have,
$E_{x}[e^{-q\tau-L(\tau)}\vartheta^{(q+m,q)}(R_{\tau},-l)\mid\mathcal{F}_{t}]$
$=\mathbf{1}\{t\leq\tau\}e^{-qt-L(t)}E_{x}[e^{-q(\tau-t)-L(\tau-t)}\vartheta^{(q+m,q)}(R_{\tau},-l)\mid\mathcal{F}_{t}]$
$\ \ \ +\mathbf{1}\{\tau<t\}e^{-q\tau-L(\tau)}\vartheta^{(q+m,q)}(R_{\tau},-l)$
$=\mathbf{1}\{t\leq\tau\}e^{-qt-L(t)}E_{R_{t}}[e^{-q\tilde{\tau}-L(\tilde{\tau})}\vartheta^{(q+m,q)}(\tilde{R}_{\tilde{\tau}},-l)]+\mathbf{1}\{\tau<t\}e^{-q\tau-L(\tau)}\vartheta^{(q+m,q)}(R_{\tau},-l)$
$=\mathbf{1}\{t\leq\tau\}e^{-qt-L(t)}\vartheta^{(q+m,q)}(R_{t},-l)+\mathbf{1}\{\tau<t\}e^{-q\tau-L(\tau)}\vartheta^{(q+m,q)}(R_{\tau},-l)$
$=e^{-q(t\wedge\tau)-L(t\wedge\tau)}\vartheta^{(q+m,q)}(R_{t\wedge\tau},-l),$
which implies the process $\{e^{-q(t\wedge\tau)-L^{(c_{1}^{*},c_{2}^{*})}(t\wedge\tau)}\vartheta^{(q+m,q)}(U^{(c_{1}^{*},c_{2}^{*})}_{t\wedge\tau},-l)\}_{t\geq 0}$ is a $P_{x}$-martingale. Alternatively, a more general verification of (3.6) can be established as follows. As derived from Proposition 2.2, the derivative of $\vartheta^{(q+m,q)}(\cdot,-l)$ fails to exist at 0 or $b$ when $X$ is of BV, and the second left derivative of $\vartheta^{(q+m,q)}(\cdot,-l)$ at $c_{2}^{*}$ does not necessarily equal zero when $X$ is of UBV. Accordingly, $(\mathcal{A}-q-m\mathbf{1}\{x<0\})V_{(c_{1}^{*},c_{2}^{*})}(x)$ is not well-defined. Therefore, we moderate our claim to demonstrate that the following result holds for any $t\in[0,T]$,
$\int_{0}^{t}e^{-qs-L^{(c_{1}^{*},c_{2}^{*})}(s)}(\mathcal{A}-q-m\mathbf{1}\{\tilde{U}_{s}^{(c_{1}^{*},c_{2}^{*})}<0\})V_{(c_{1}^{*},c_{2}^{*})}(\tilde{U}_{s}^{(c_{1}^{*},c_{2}^{*})})\mathrm{d}s\leq 0,\ \ \ a.s.,$ (4.13)
@@ 5 Examples
The existence and satisfiability of the optimal impulse control $\pi_{(c_{1}^{*},c_{2}^{*})}$ are showed in Proposition 4.2. Combined with Theorem 4.6, this result enables the development of a method for numerically computing the optimal thresholds. Compared to the de Finetti’s problem discussed in Loeffen[26], the additional complexity involves characterizing the optimal values of parameters $c_{1}$ and $c_{2}$, which forms a two-dimensional optimization problem, in contrast to the one-dimensional problem of finding the optimal barrier strategy. In this section, we analyze two examples for which it is able to present the numerical results of the optimal points $(c_{1}^{*},c_{2}^{*})$, including refracted Brownian risk model and refracted Cramér-Lundberg model with exponential claims.